OpenStudy (anonymous):

what does 5√3 means?

OpenStudy (saifoo.khan):

5 square root 3.

OpenStudy (turingtest):

That question seems philosophical...

OpenStudy (anonymous):

it could be the 5th root of 3 too !

OpenStudy (turingtest):

true true!

OpenStudy (anonymous):

√ = square root therefore it means 5 square root 3

OpenStudy (saifoo.khan):

It couldnt be fifth root of 3.

OpenStudy (anonymous):

no that symbol only means root !

OpenStudy (turingtest):

yes but it is written quite wring technically, and could be interpreted either way people write x2 for x^2 here all the time.

OpenStudy (anonymous):

The sum of the first n consecutive odd whole numbers i y, where n>0. √y-n =?

OpenStudy (anonymous):

is y*

OpenStudy (turingtest):

I really don't understand the question now. You mean a sum with a sigma sign?

OpenStudy (anonymous):

root(y - n) = (n-1)root(n)

OpenStudy (anonymous):

fifth root of 3 is ====> 1.24573094

OpenStudy (anonymous):

\[\sqrt{y-n} - (n-1)\sqrt{n}\]

OpenStudy (turingtest):

minus or equals? and what are you doing, can you explain at all?

OpenStudy (turingtest):

@harkirat

OpenStudy (anonymous):

Yes, i will. Just wait a few minutes.... I am preparing it......

OpenStudy (anonymous):

The sum of the first n consecutive odd whole numbers i y, where n>0. √y-n =? take first two odd whole numbers --- 1 and 3 so n=2 y = 1+ 3 = 4 y - n = 4-2 = 2 root(y-n) = root(2) = (2-1)root(4) i.e (n-1)root(n)

OpenStudy (anonymous):

take first five consecutive odd whole numbers --- 1, 3, 5, 7 and 9 so n=5 y = 1+3+5+7+9 = 25 y - n = 25-5 = 20 root(y-n) = root(20) = 4 root(5) = (5-1)root(5) i.e (n-1)root(n)

OpenStudy (anonymous):

Is the logic clear now...????

OpenStudy (anonymous):

Sorry I made a mistake in the equation !!!!\[\sqrt{y-n} = (n - 1)\sqrt{n}\]

OpenStudy (anonymous):

but the choices are... A. n B.-n C.-y D.0 E. It cannot be determined by the information givin

OpenStudy (anonymous):

In your question is it Root(y - n) or root(y) - n ????

OpenStudy (anonymous):

\[\sqrt{y - n} or \sqrt{y} - n\]

OpenStudy (anonymous):

Question hai.. The sum of the first n consecutive odd whole numbers is y, where n>0. √y-n

OpenStudy (anonymous):

question is.. The sum of the first n consecutive odd whole numbers is y, where n>0. √y-n

OpenStudy (turingtest):

@harkirat, that's very interesting, thanks!

OpenStudy (anonymous):

@mathshasproblems Are u from india ?? I guess because you used the word "hai"..... I'm also from India (Delhi)..

OpenStudy (anonymous):

@TuringTest Just struck me.......after I wrote it on paper and tried to find a pattern....

OpenStudy (anonymous):

It could also be \[\sqrt{28}\]

OpenStudy (anonymous):

Got the answer !!!!!

OpenStudy (anonymous):

The answer is D. 0

OpenStudy (anonymous):

I think we have to find \[\sqrt{y}-n = ?\] The sum of the first n consecutive odd whole numbers i y, where n>0. √y-n =? take first two odd whole numbers --- 1 and 3 so n=2 y = 1+ 3 = 4 root(y) = root(4) = 2 so root(y) - n = 2-2 = 0 take first five consecutive odd whole numbers --- 1, 3, 5, 7 and 9 so n=5 y = 1+3+5+7+9 = 25 root(y) = root(25) = 5 so root(y) - n = 5-5 = 0

OpenStudy (turingtest):

Nice, though the other problem seemed harder. Hat's off to you Harkirat.

OpenStudy (anonymous):

@TuringTest Thanks for yr nice comment..☺☺ Once he sent the answers, I just had to check again and yes it is very simple this time!!